Sarah is 3 times as old as Rob. their ages together add up to 16. how old is Sarah?
Answer by ewatrrr(24785)

Hi,
Let x be Rob's age. Then 3x will be Sarah's age.
.
question states
x + 3x = 16
4x = 16
x = 4 (Rob's age)
.
3x = 12
Sarah is 12 years old.
.
check your answer
4 + 12 = 16